Find the general/particular solution of the following Differential Equations

(Exact D.E)



3.) [e^x + lny+(y/x)]dx [(x/y)+lnx+siny]dy=0


1
Expert's answer
2021-09-07T02:51:42-0400

Solution;

Check for exactiness of the equation;

"M=e^x+lny+\frac yx"

"N=\frac xy+lnx+siny"

"\frac{dM}{dy}=\frac 1y+\frac 1x"

"\frac{dN}{dx}=\frac 1y+\frac 1x"

Clearly;

"\frac{dM}{dy}=\frac{dN}{dx}"

The equation is exact.

Solution is given by;

"\int_{y=c}Mdx+\int" (terms in N independent of x)"dy=C"

"\int(e^x+lny+\frac yx)dx+\int sin (y )dy=C"

"\int e^xdx+lny\int1dx+y\int\frac 1xdx+\int sin (y)dy=C"

Solve to obtain the general solution;

"e^x+xln(y)+yln(x)-cos(y)=C"